# Assertion-Conditioned Compliance: A Provenance-Aware Vulnerability in Multi-Turn Tool-Calling Agents

### v1

@@ 1 Introduction
Large Language Models (LLMs) augmented with the ability to invoke external tools (tool-calling) have demonstrated remarkable capabilities beyond their standalone performance li-etal-2023-api; qu2024toollearninglargelanguage; wang2024toolsanywaysurveylanguage. However, their successful deployment in safety-critical industries remains constrained by concerns about transparency and the evaluation of their reasoning robustness Liao2024AI. When such agents encounter incorrect or contradictory information — an event familiar in real-world user interactions with LLMs feng-etal-2025-unraveling, and a serious concern in unmaintained deployments — an agent that treats this feedback as authoritative can propagate incorrect states and cause harmful downstream degradation in both the tool-calling pipeline and the real-world systems it influences yao2023reactsynergizingreasoningacting.
@@
Previous work has documented linguistic sycophancy sharma2025understandingsycophancylanguagemodels; cheng2025elephantmeasuringunderstandingsocial and has confirmed that tool-augmented agents are vulnerable to safety issues arising from both misleading user inputs and contradictory tool data yeetal2024toolsword. However, the provenance of these misleading signals—and how an agent’s compliance bias differs when the source is the User (a social cue) versus the Function (a system cue)—remains poorly understood and has yet to be systematically evaluated.
@@
Figure 1: Overview of baseline, user-sourced assertion (USA), and function-sourced assertion (FSA) behavior in multi-turn tool-calling. Assertions may cause compliance in the immediate function call, after which the model may either propagate the misleading path (degradation) or recover and produce the correct final environment state.
@@
We introduce Assertion-Conditioned Compliance (A-CC) to formally diagnose this gap. A-CC is a failure mode in which model accepts an incorrect assertion and updates its internal reasoning to incorporate it, thus manipulating its execution pipeline. This goes beyond standard linguistic sycophancy, which is limited to verbal agreement in the surface response. A-CC instead captures procedural compliance, in which the execution pipeline itself is altered. In this setup, provenance is central. We measure compliance across two distinct vectors: plausible but incorrect user-sourced assertions (USAs), which test sycophancy’s effect on tool-use, and contradictory function-sourced assertions (FSAs), which test whether the agent allows erroneous tool feedback to redirect its execution pipeline. This latter FSA vector models a plausible and insidious industrial risk (where, for example, stale hints may originate from unmaintained tools in mismanaged deployments; while unlikely, this would have a global effect across said deployments) (illustrated in Figure 1).
@@
We ground our evaluation in the BFCL patil2025the, extending its multi-turn, state-dependent tasks from a simple accuracy metric to a diagnostic of procedural robustness in realistic tool APIs. Our experiments show how A-CC exposes a consistent behavioral weakness across model families and sizes: plausible assertions reliably reshape tool-calling pipelines, even when final BFCL accuracy remains high. Importantly, we find that assertion compliance is not tightly coupled to accuracy degradation, indicating that assertion-following carries risks beyond the degradation of the user’s task. For industry settings, this means that models judged solely on task success accuracy by leaderboard scores may still execute unnecessary or unsafe operations under natural user prompt variation or stale tool hints. By making provenance and procedural compliance explicit, A-CC provides a practical diagnostic that reinforces the need for additional guardrails and safety measures. In summary, our contributions are as follows:
@@
  • A compliance-based evaluation metric (compliance rate) that decouples obedience to asserted operations from task success, quantifying how often agents execute unnecessary or unsafe functions even when pipelines are otherwise correct.
@@ 2 Related Work
Sycophancy in Large LMs. Prior work has shown that LLMs trained via RLHF often produce outputs that align with a user’s expressed belief or preferences, even when doing so sacrifices accuracy or validity sharma2025understandingsycophancylanguagemodels. wei2024simplesyntheticdatareduces extend this line of work by quantifying sycophancy across opinion and arithmetic-based settings, showing that a lightweight synthetic-data intervention can reduce it. More recently, liu2025truthdecayquantifyingmultiturn move beyond single-turn prompts and introduce a benchmark for evaluating multi-turn sycophancy, demonstrating that conformity to user beliefs can intensify over extended dialogue. Our A-CC framework is complementary: rather than focusing on agreement in surface text, we study how such assertion-following manifests into procedural compliance in the function-calling pipeline of tool-augmented agents.
@@
Benchmarks for function-calling. Our study builds on the multi-turn tasks of the BFCL patil2025the, using its real-world APIs and final-environment-state scoring as a natural substrate for analyzing assertion-conditioned robustness. BFCL’s multi-turn category couples stateful tool interactions with a fixed success metric, which allows us to isolate how provenance-tagged assertions reshape the execution pipeline without modifying the underlying benchmark. Other function-calling benchmarks are complementary but target different aspects of the problem.
@@
Robustness and adversarial contexts in tool use. A growing strand of research examines failures when the “environment around” the agent shifts or is hostile. Prior work shows that naturalistic prompt variation, toolkit expansion with semantically similar tools, and distributional shifts in tool selection all induce agentic degradation and can collapse certified worst-case accuracy under adversarial tool perturbations (rabinovichanabytavor2025robustness; yeon2025quantifyingdistributionalrobustnessagentic). Other work also demonstrates how deceptive or malicious tool injections can trigger privacy leaks, DoS actions, or unintended executions, and that tool-calling modes expose jailbreak paths not visible in standard chat interfaces (zhangetal2025allies). Stage-based safety suites such as ToolSword yeetal2024toolsword stress-tests input, execution, and output stages of tool-calling LLMs under explicitly specified safety scenarios, measuring attack success, unsafe tool selection, and harmful or erroneous feedback propagation; our work is complementary in that we study how non-adversarial, provenance-tagged assertions within standard benchmarks reshape procedural compliance, rather than curating new safety scenarios.
@@
Alignment and decision-making in tool use. Beyond defenses, alignment-oriented work specifies desirable behaviors. chenetal2024towardstool proposes H2A—Helpfulness, Harmlessness, Autonomy—and releases ToolAlign to train models that (i) call tools to help, (ii) refuse unsafe instructions and insecure tool responses, and (iii) avoid unnecessary calls. This closely matches A-CC’s desiderata (non-compliance with harmful or misleading assertions, restraint in over-calling).
@@
Capability expansion and training for function calling. Orthogonal to robustness, several efforts have improved coverage or accuracy. qinetal2025meta’s Meta-Tool retrieves appropriate tools from large libraries (and introduces Meta-Bench), thereby boosting open-world tool selection and enabling smaller models to rival larger baselines. This approach is helpful for breadth, but not designed to resist misinformation in context. chenetal2025enhancing report prompt and data strategies—including instruction mixtures, “Decision Tokens,” and multilingual tool pipelines—that improve tool-choice relevance and overall function-call accuracy. Our A-CC evaluation instead asks a complementary question: given existing tool capabilities, how stable is the agent’s decision-making pipeline when confronted with plausible but incorrect assertions, and how often does compliance with such assertions translate into unnecessary or harmful tool executions?
@@ 3.1 Benchmark & Task Setup (BFCL)
Our experiment is built and evaluated using the BFCL patil2025the. Task success is determined by the final environment state, not merely the sequence of tools invoked, allowing for variations in reasoning across the entire pipeline. We exclusively use the multi_turn_base category, a set of 200 curated entries inspired by common real-world APIs.
@@ 3.2.1 User-Sourced Assertions (USAs)
Assertions in our framework are characterized by provenance — the source of the misleading signal in the multi-turn pipeline. User-sourced assertions (USAs) represent social provenance, simulating sycophancy by providing misleading information from the user prompt itself. Our generation of USAs simulates user-level misdirection by constructing plausible but incorrect claims about which function-call is appropriate. Generation inputs include the original BFCL prompt, documentation snippets, and a set of incorrect function candidates (derived from BFCL documentation). Each assertion is a single sentence authored under two tonal variants — confident and hedged — and three functional modes: init, read-heavy, and write-heavy. These modes target specific operational contexts: read-heavy and write-heavy assertions are designed for turns with a high density of available read or write functions, respectively, and assert a plausible but incorrect function of that type. At evaluation time, these generated assertions are injected into the corresponding BFCL turns according to our data injection and execution protocol (cf.§3.3).
This targeting allows us to test distinct behavioral contexts: read-heavy turns involve information-retrieval functions and tend to have more subtle implications, such as unnecessarily surfacing additional or sensitive information. Write-heavy turns involve environment-mutating functions and thus revealing the most severe degradations. Init represents a mixed boundary condition combining both read and write patterns. Each generated record includes metadata specifying the target function, assertion text, and turn ID, which allows reproducible analysis of the effects of linguistic compliance.
@@ 3.2.2 Function-Sourced Assertions (FSAs)
In contrast, function-sourced assertions (FSAs) capture procedural provenance — cases in which misleading procedural suggestions originate from the agent’s internal tool response rather than the user. This isolates how agents treat conflicting or stale system feedback as authoritative signals, offering insight into the model’s internal authority bias. Each FSA is injected into a specific tool’s output as a system policy note (a textual note appended to the tool’s response payload, similar in implementation to lu2025dontjustfinetuneagent), declaring a rule (within the same tool category) contradicting the user’s intended actions.
Our experiment structures the FSA set as an ablation condition targeting only write-heavy turns/functions in confident language (given the limited variation a policy hint would have in its tone or implementation for more critical functions) to test the agent’s compliance hierarchy. Concretely, for each selected write-heavy turn we inject a single operator hint into the tool output in the absence of any USA; this corresponds to the “FSA Baseline” reported in our results. Each generated record also includes the aforementioned metadata but includes a (turn_idx, func) flag to indicate where the assertion would be injected during evaluation.
@@ 3.3 Data Injection & Execution Protocol
  • For USAs, we modify the BFCL input JSON by replacing the targeted user turn with the asserted sentence, preserving the original turn ID.
  • For FSAs, we inject the operator-level hint directly into the tool’s function output. To mirror realistic "misleading" tool feedback, this injection is conditional: it only occurs if the agent successfully invokes the target function at the intended turn.
@@ 3.4 Metrics
Model
No-assert
succ. (%)
USA comp. (CR) FSA comp. (CR)
Worst
$\Delta$ succ.
Conf. Hedg. Conf. Hedg.
BitAgent Bounty 8B 77.7 33.3 21.3 21.8 20.3 -20.8
Qwen3 32B (FC) 54.3 34.5 26.2 40.6 19.8 -19.3
Qwen3 14B (FC) 50.2 33. 26.2 41.6 19.3 -23.4
Qwen3 8B (FC) 43.1 30.6 22.3 31.5 17.8 -14.7
xLAM 2 70B FC r 79.2 38.6 29.6 28.4 25.9 -17.8
xLAM 2 32B FC r 80.7 34.7 27.3 26.4 18.8 -20.3
xLAM 2 8B FC r 77.2 33.5 22.7 34.0 20.3 -11.7
xLAM 2 3B FC r 70.6 28.4 21.3 24.9 14.7 -14.2
Watt Tool 70B 70.0 47.5 37.6 37.6 32.5 -16.8
Watt Tool 8B 45.2 37.7 28.6 36.5 27.9 -10.2
ToolACE 2 8B 46.7 47.2 41.3 21.8 25.9 -16.8
Table 1: Summary of baseline BFCL multi-turn success (No-assert succ.), assertion-conditioned compliance rates (CR) under user-sourced and function-sourced assertions (USA / FSA, confident vs. hedged), and worst-case success degradation (Worst $\Delta$ succ.) over (non-interaction) A-CC conditions for the top 11 models on the BFCL leaderboard. Baseline success ranges from 43.1% (Qwen3 8B (FC)) to 80.7% (xLAM 2 32B FC r), with a macro-average of 63.2%. Larger xLAM and Watt Tool variants dominate the no-assert baseline (around 79–81%), while smaller Qwen3 and ToolACE models consistently lag behind. Derived from Table 3 in the Appendix.
Our evaluation relies on two complementary metrics: task success (accuracy) and compliance rate (CR). A central goal of our work is to analyze their relationship —specifically, to test whether task accuracy and behavioral compliance are not closely correlated. We posit that accuracy degradation alone is insufficient to capture the behavioral risks of assertion-following, as agents can often comply with dangerous assertions while still achieving task success (i.e. behavioral compliance can have inconspicuous impacts on the execution environment regardless of the pipeline result). Note that task success represents the standard BFCL accuracy score (and, as per the benchmark’s rules, success is determined by the final environment state at the end of the multi-turn dialogue).
We introduce compliance rate (CR) as our primary behavioral metric. CR captures whether the model incorporates the asserted operation by invoking the asserted function at least once at any point within the same turn the assertion is made. Each evaluation case is determined to either be compliant (the asserted operation is invoked) or non-compliant (the model never invokes it). This metric isolates behavioral adoption from task success, allowing us to quantify how often an assertion actually alters the agent’s decision-making process, even in cases where the final BFCL score is correct.
To further analyze the relationship between accuracy and compliance, we also pair each asserted run with its corresponding no-assertion baseline and compute success transitions into 4 distinct "outcome buckets":
  • S $\to$ S (success preserved)
  • S $\to$ F (assertion-induced failure)
  • F $\to$ S (assertion-induced recovery - unlikely)
  • F $\to$ F (failure persists)

### v2

@@ 1 Introduction
Large Language Models (LLMs) augmented with the ability to invoke external tools (tool-calling) have demonstrated remarkable capabilities beyond their standalone performance Li et al. (2023); Qu et al. (2024); Wang et al. (2024). However, their successful deployment in safety-critical industries remains constrained by concerns about transparency and the evaluation of their reasoning robustness Liao and Wortman Vaughan (2024). When such agents encounter incorrect or contradictory information — an event familiar in real-world user interactions with LLMs Feng et al. (2025), and a serious concern in unmaintained deployments — an agent that treats this feedback as authoritative can propagate incorrect states and cause harmful downstream degradation in both the tool-calling pipeline and the real-world systems it influences Yao et al. (2023).
@@
Previous work has documented linguistic sycophancy Sharma et al. (2025); Cheng et al. (2025) and has confirmed that tool-augmented agents are vulnerable to safety issues arising from both misleading user inputs and contradictory tool data Ye et al. (2024). However, the provenance of these misleading signals — and how an agent’s compliance bias differs when the source is the User (a social cue) versus the Function (a system cue) — remains poorly understood and has yet to be systematically evaluated.
@@
Figure 1: Overview of baseline, user-sourced assertion (USA), and function-sourced assertion (FSA) behavior in multi-turn tool-calling. Assertions may cause compliance in the immediate function call, after which the model may either propagate the misleading path (degradation) or recover and produce the correct final environment state. The bottom panels illustrates the resulting risk scope: USAs pose a localized, direct risk to the immediate session, whereas FSAs introduce broad, indirect risks by potentially polluting the shared environment for multiple users.
@@
We introduce Assertion-Conditioned Compliance (A-CC) to formally diagnose this gap. A-CC is a failure mode in which the model accepts an incorrect assertion and updates its internal reasoning to incorporate it, thus manipulating its execution pipeline. This goes beyond standard linguistic sycophancy, which is limited to verbal agreement in the surface response. A-CC instead captures procedural compliance, in which the execution pipeline itself is altered. In this setup, provenance is central. We measure compliance across two distinct vectors: plausible but incorrect user-sourced assertions (USAs), which test sycophancy’s effect on tool-use, and contradictory function-sourced assertions (FSAs), which test whether the agent allows erroneous tool feedback to redirect its execution pipeline. This latter FSA vector models a plausible and insidious industrial risk (where, for example, stale hints may originate from unmaintained tools in mismanaged deployments; while unlikely, this would have a global effect across said deployments) (illustrated in Figure 1).
We distinguish our methodology from standard prompt injection or tool misuse, which typically aim to hijack the agent’s goal for malicious purposes (bypassing safety filters, breaching databases, etc.). In contrast, A-CC demonstrates contextual trust: the agent maintains its helpful objective but fails to verify false premises against the tool history, environment state or function documentation. This represents a failure of grounding rather than alignment, as the model prioritizes user/function-sourced hallucinations over execution reality.
@@
We ground our evaluation in the BFCL Mao et al. (2024), extending its multi-turn, state-dependent tasks from a simple accuracy metric to a diagnostic of procedural robustness in realistic tool APIs. Our experiments show how A-CC exposes a consistent behavioral weakness across model families and sizes: plausible assertions reliably reshape tool-calling pipelines, even when final BFCL accuracy remains high. Importantly, we find that assertion compliance is not tightly coupled to accuracy degradation, indicating that assertion-following carries risks beyond the degradation of the user’s task. For industry settings, this means that models judged solely on task success accuracy by leaderboard scores may still execute unnecessary or unsafe operations under natural user prompt variation or stale tool hints. By making provenance and procedural compliance explicit, A-CC provides a practical lens that reinforces the need for additional guardrails and safety measures. In summary, our contributions are as follows:
@@
  • A compliance-based evaluation metric (compliance rate) that decouples obedience (to asserted operations) from task success, quantifying how often agents execute unnecessary or unsafe functions even when pipelines are evaluated to be correct.
@@ 2 Related Work
Sycophancy in Large LMs. Prior work has shown that LLMs trained via RLHF often produce outputs that align with a user’s expressed belief or preferences, even when doing so sacrifices accuracy or validity Sharma et al. (2025). Wei et al. (2024) extend this line of work by quantifying sycophancy across opinion and arithmetic-based settings, showing that a lightweight synthetic-data intervention can reduce it. More recently, Liu et al. (2025a) move beyond single-turn prompts and introduce a benchmark for evaluating multi-turn sycophancy, demonstrating that conformity to user beliefs can intensify over extended dialogue. Our A-CC framework is complementary: rather than focusing on agreement in surface text, we study how such assertion-following manifests into procedural compliance in the function-calling pipeline of tool-augmented agents.
@@
Benchmarks for function-calling. Our study builds on the multi-turn tasks of the BFCL Mao et al. (2024), using its realistic API definitions and final-environment-state scoring as a natural substrate for analyzing assertion-conditioned robustness. BFCL’s multi-turn category couples stateful tool interactions with a fixed success metric, which allows us to isolate how provenance-tagged assertions reshape the execution pipeline without modifying the underlying benchmark.
@@
Robustness and adversarial contexts in tool use. A growing strand of research examines failures when the “environment around” the agent shifts or is hostile. Prior work shows that naturalistic prompt variation, toolkit expansion with semantically similar tools, and distributional shifts in tool selection all induce agentic degradation and can collapse certified worst-case accuracy under adversarial tool perturbations (Rabinovich and Anaby Tavor, 2025; Yeon et al., 2025). Other work also demonstrates how deceptive or malicious tool injections can trigger privacy leaks, DoS actions, or unintended executions, and that tool-calling modes expose jailbreak paths not visible in standard chat interfaces (Zhang et al., 2025b). Stage-based safety suites such as ToolSword Ye et al. (2024) stress-tests input, execution, and output stages of tool-calling LLMs under explicitly specified safety scenarios, measuring attack success, unsafe tool selection, and harmful or erroneous feedback propagation; our work is complementary in that we study how non-malicious, provenance-tagged assertions within standard benchmarks reshape procedural compliance, rather than curating new safety scenarios.
@@
Alignment and decision-making in tool use. Beyond defenses, alignment-oriented work specifies desirable behaviors. Chen et al. (2024) proposes H2A — Helpfulness, Harmlessness, Autonomy — and releases ToolAlign to train models that (i) call tools to help, (ii) refuse unsafe instructions and insecure tool responses, and (iii) avoid unnecessary calls. This closely matches A-CC’s desiderata (non-compliance with harmful or misleading assertions, restraint in over-calling).
@@
Capability expansion and training for function calling. Orthogonal to robustness, several efforts have improved coverage or accuracy. Qin et al. (2025)’s Meta-Tool retrieves appropriate tools from large libraries (and introduces Meta-Bench), thereby boosting open-world tool selection and enabling smaller models to rival larger baselines. This approach is helpful for breadth, but not designed to resist misinformation in context. Chen et al. (2025) report prompt and data strategies — including instruction mixtures, “Decision Tokens,” and multilingual tool pipelines — that improve tool-choice relevance and overall function-call accuracy. Our A-CC evaluation instead asks a complementary question: given existing tool capabilities, how stable is the agent’s decision-making pipeline when confronted with plausible but incorrect assertions, and how often does compliance with such assertions translate into unnecessary or harmful tool executions?
@@ 3.1 Benchmark & Task Setup (BFCL)
Our experiment is built and evaluated using the BFCL Mao et al. (2024). Task success is determined by the final environment state, not merely the sequence of tools invoked, allowing for variations in reasoning across the entire pipeline. We exclusively use the multi_turn_base category, a set of 200 curated entries inspired by common real-world APIs.
@@ 3.2.1 User-Sourced Assertions (USAs)
Assertions in our framework are characterized by provenance — the source of the misleading signal in the multi-turn pipeline. User-sourced assertions (USAs) represent social provenance, simulating sycophancy by providing misleading information from the user prompt itself. Our generation of USAs simulates user-level misdirection by constructing plausible but incorrect claims about which function-call is appropriate. Generation inputs include the original BFCL prompt, documentation snippets, and a set of incorrect function candidates (derived from BFCL documentation). Each assertion is a single sentence authored under two tonal variants — confident and hedged — and three functional modes: init, read-heavy, and write-heavy. These modes target specific operational contexts: read and write-heavy assertions are designed for turns with a high density of available read or write functions, respectively, and assert a plausible but incorrect function of that type. At evaluation time, these generated assertions are injected into the corresponding BFCL turns according to our data injection and execution protocol (cf.§3.3).
This targeting allows us to test distinct behavioral contexts: read-heavy turns involve information-retrieval functions and tend to have more subtle implications, such as unnecessarily surfacing additional or sensitive information. Write-heavy turns involve environment-mutating functions and thus reveals the most severe degradations. Init represents a mixed boundary condition combining both read and write patterns. Each generated record includes metadata specifying the target function, assertion text, and turn ID, which allows reproducible analysis of the effects of linguistic compliance.
@@ 3.2.2 Function-Sourced Assertions (FSAs)
In contrast, function-sourced assertions (FSAs) capture procedural provenance — cases in which misleading procedural suggestions originate from the agent’s internal tool response rather than the user. This isolates how agents treat conflicting or stale system feedback as authoritative signals, offering insight into the model’s internal authority bias. Each FSA is injected into a specific tool’s output as a system policy note (a textual note appended to the tool’s response payload, similar in implementation to Lu et al. (2025)), declaring a rule (within the same tool category) contradicting the user’s intended actions.
Our experiment structures the FSA set as an ablation condition targeting only write-heavy turns/functions in confident language (given the limited variation a policy hint would have in its tone or implementation for more critical functions) to test the agent’s compliance hierarchy. Concretely, for each selected write-heavy turn we inject a single operator hint into the tool output in the absence of any USA; this corresponds to the “FSA Baseline” reported in our results. Each generated record also includes the aforementioned metadata and a (turn_idx, func) flag to indicate where the assertion would be injected during evaluation.
3.2.3 Data Validation
To ensure the validity of our synthetic assertions, we employed a human-in-the-loop verification protocol across all of our generated datasets. Authors assessed samples for semantic validity, tonal consistency (confident vs. hedged), and contextual plausibility. To maintain statistical rigor, any batch containing unnatural artifacts or logic errors was discarded in its entirety and regenerated (so as to avoid potential “authorial bias” from manual modifications). This process ensures that the distribution of assertions remains controlled while satisfying the naturalness requirements of realistic user-agent interactions.
@@ 3.3 Data Injection & Execution Protocol
Model No-assert succ. USA comp. (CR) FSA comp. (CR) Worst $\Delta$ succ.
Conf. Hedg. Abl. set baseline
BitAgent Bounty 8B 77.7 33.3 21.3 18.3 -20.8
Qwen3 32B (FC) 54.3 34.5 26.2 41.1 -19.3
Qwen3 14B (FC) 50.2 33.0 26.2 40.1 -23.4
Qwen3 8B (FC) 43.1 30.6 22.3 27.9 -14.7
xLAM 2 70B FC r 79.2 38.6 29.6 22.3 -17.8
xLAM 2 32B FC r 80.7 34.7 27.3 29.9 -20.3
xLAM 2 8B FC r 77.2 33.5 22.7 21.8 -11.7
xLAM 2 3B FC r 70.6 28.4 21.3 23.4 -14.2
Watt Tool 70B 70.0 47.5 37.6 32.0 -16.8
Watt Tool 8B 45.2 37.7 28.6 17.3 -10.2
ToolACE 2 8B 46.7 47.2 41.3 32.0 -16.8
Table 1: Performance summary for the top 11 models on the BFCL leaderboard. We report baseline multi-turn success (No-assert succ.) alongside three A-CC metrics: (1) USA compliance rates (CR), macro-averaged across init/read-heavy/write-heavy sets; (2) FSA compliance rates, calculated on the ablation set; and (3) Worst-case success degradation (Worst $\Delta$ succ.) across non-interaction conditions. No-assert success ranges from 43.1% (Qwen3 8B (FC)) to 80.7% (xLAM 2 32B FC r), with a macro-average of 63.2%. Larger xLAM and Watt Tool variants dominate the baseline ($\sim$80%), while smaller Qwen3 and ToolACE models consistently lag behind. Derived from Tables 3, 4, 5 & 6 in the Appendix. Total $n=197$ cases.
@@
  • For USAs, we modify the BFCL input JSON by injecting the targeted user turn with the asserted sentence (preserving the original turn ID).
  • For FSAs, we inject the operator-level hint directly into the tool’s function output. To mirror realistic “misleading” tool feedback, this injection is conditional: it only occurs if the agent successfully invokes the target function at the intended turn.
@@ 3.4 Metrics
Our evaluation relies on two complementary metrics: task success (accuracy) and compliance rate (CR). A central goal of our work is to analyze their relationship — specifically, to test whether task accuracy and behavioral compliance are not closely correlated. We posit that accuracy degradation alone is insufficient to capture the behavioral risks of assertion-following, as agents can often comply with dangerous assertions while still achieving task success (i.e., behavioral compliance can have inconspicuous impacts on the execution environment regardless of the pipeline result). Note that task success represents the standard BFCL accuracy score (and, as per the benchmark’s rules, success is determined by the final environment state at the end of the multi-turn dialogue).
We utilize compliance rate (CR) as our primary behavioral metric. CR captures whether the model incorporates the asserted operation by invoking the asserted function at least once at any point within the same turn the assertion is made. Each evaluation case is determined to either be compliant (the asserted operation is invoked) or non-compliant (the model never invokes it). This metric isolates behavioral adoption from task success, allowing us to quantify how often an assertion actually alters the agent’s decision-making process, even in cases where the final evaluation is deemed correct.
To further analyze the relationship between accuracy and compliance, we also pair each asserted run with its corresponding no-assertion baseline and compute success transitions into 4 distinct “outcome buckets”:
[…]